OpenStudy (anonymous):

Solve x2 – 2x = 8 by completing the square. Show your work for full credit.

OpenStudy (anonymous):

\[x ^{2}-2x-8=0\] (x+2)(x-4)=0 x+2=0 x-4=0 x=-2 & x=4

OpenStudy (anonymous):

x2 – 2x +1 -1= 8 (x-1)2 -1 =8 (x-1)2 =9 x-1=+-3 x=4 x=-2

OpenStudy (anonymous):

my way is what you looking for @dmartinez130. @katlg87 solved it by factorizing

OpenStudy (anonymous):

\[x^{2}-2x=8\] \[x^{2}-2x-8=0\] \[x^{2}-4x+2x-8=0\] x(x-4)+2(x-4)=0 (x+2) (x+4) x=-2 OR x=4

OpenStudy (campbell_st):

rewrite the equation \[x2 -2x = 8\] the general form is \[ax^2 + bx + c\] to complete the square add \[(\frac{b}{2})^2\] to both sides in you question add (-2/2)^1 to both sides \[x^2 -2x + 1 = 9\] now factorise the left hand side \[(x -1)^2 = 9 \]